# On $3$-manifolds admitting co-orientable taut foliations, but none with vanishing Euler class

### v2

@@ 5.3. $\mathcal{B}$ is hyperbolic
(2) For $n\geq 3$, let $M=M(0;2-n,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})$, where $\sum_{i}\frac{a_{i}}{b_{i}}\leq\frac{n-2}{2}$. Then $M$ has a horizontal foliation, but $e(\nu_{M})=0$ if and only if $a_{i}=1$ for $1\leq i\leq n$.
(3) For $n\geq 4$, consider the manifold $M=M(0;b,\frac{1}{b_{1}},\ldots,\frac{1}{b_{n}})$, where the $b_{i}$ are pairwise coprime and $\geq n$. Then
  • (a)
    $M$ has a horizontal foliation if and only if $2-n\leq b\leq-1$.
  • (b)
    $e(\nu_{M})=0$ if and only if b=2nb=2-n.
(4) For $n\geq 4$, let $M=M(0;b,\frac{a_{1}}{b_{1}},\ldots,\frac{a_{n}}{b_{n}})$, where $\sum_{i}\frac{a_{i}}{b_{i}}\leq\frac{n}{6}$. Then
  • (a)
    $M$ has a horizontal foliation for $2-n\leq b\leq-2$, and for b=1b=-1 if $n=4,5$, or $6$.
  • (b)
    $e(\nu_{M})=0$ for at most one $b$ with $2-n\leq b\leq\frac{2-n}{2}$.

### v3

@@ 5.3. $\mathcal{B}$ is hyperbolic
    $M$ has a horizontal foliation if and only if up to orientation, i.e. for either $M$ or M-M, $a_{1}=\ldots=a_{n}=1$ and b=2nb=2-n. Then taking $m=1$ shows that $e(\nu_{M})=0$.
  2. (2)
    There exists $M$ with a horizontal foliation such that for both $M$ and M-M either $b\neq 2-n$ or some $a_{i}\neq 1$, but any such $M$ has $e(\nu_{M})=0$.
  3. (3)
    There exists $M$ with a horizontal foliation having $e(\nu_{M})\neq 0$, but any such $M$ has $e(M)=0$, and hence is not a rational homology sphere.
The exceptions of Type (1) can be easily determined using Remark 5.4. They are
  • $S^{2}(2,3,c)$, where $c=7,8,9,10,12,14,18,24,$ or $30$
  • $S^{2}(2,4,c)$, where $c=5,6,8,$ or $12$
  • $S^{2}(2,5,c)$, where $c=5$ or 66
  • $S^{2}(2,6,c)$, where $c=6$ or $12$
  • $S^{2}(3,3,c)$, where $c=4$ or 66
  • $S^{2}(3,4,c)$, where $c=4$ or 66
  • $S^{2}(4,4,c)$, where $c=4$ or 66
  • $S^{2}(6,6,6)$
  • $S^{2}(2,2,2,c)$, where $c=3,4,$ or 66
  • $S^{2}(2,2,2,2,2)$
For the exceptions of Type (2), we list the corresponding manifolds $M$ up to orientation, and the values of $m$ which show that $e(\nu_{M})=0$.
  • $S^{2}(2,3,c)$, where $c=11,13,16$, or $20$ and M=M(0;1,1/2,1/3,a/c)M=M(0;-1,1/2,1/3,a/c), $1<a<c/2$. In particular,
    • when $c=11$, $a=2$ and m=5m=-5
    • when $c=13$, $a=2$ and $m=7$
    • when $c=16$, $a=3$ and m=5m=-5
    • when $c=20$, $a=3$ and $m=7$
  • $S^{2}(2,4,c)$, where $c=7,9,10,14,18,$ or $30$ and M=M(0;1,1/2,1/4,a/c)M=M(0;-1,1/2,1/4,a/c), $1<a<c/2$. In particular,
    • when $c=7$, $a=2$ and m=3m=-3
    • when $c=9$, $a=2$ and $m=5$
    • when $c=10$, $a=3$ and m=3m=-3
    • when $c=14$, $a=3$ and $m=5$
    • when $c=18$, $a=3$ and m=7m=-7
    • when $c=30$, $a=7$ and $m=13$
  • $S^{2}(2,6,c)$, where $c=8$ or $10$ and M=M(0;1,1/2,1/6,a/c)M=M(0;-1,1/2,1/6,a/c), $1<a<c/2$. In particular,
    • when $c=8$, $a=3$ and m=5m=-5
    • when $c=10$, $a=3$ and $m=7$
  • $S^{2}(3,3,c)$, where $c=5,7,8,$ or $10$ and M=M(0;1,1/3,1/3,a/c)M=M(0;-1,1/3,1/3,a/c), $1<a<c/2$. In particular,
    • when $c=5$, $a=2$ and m=2m=-2
    • when $c=7$, $a=2$ and m=5m=-5; or $a=3$ and m=2m=-2
    • when $c=8$, $a=3$ and m=5m=-5
    • when $c=10$, $a=3$ and $m=7$
  • $S^{2}(3,4,5)$, where M=M(0;1,1/3,1/4,2/5)M=M(0;-1,1/3,1/4,2/5) and $m=13$
  • $S^{2}(4,4,12)$, where M=M(0;1,1/4,1/4,5/12)M=M(0;-1,1/4,1/4,5/12) and $m=5$
  • $S^{2}(4,6,6)$, where $M=M(3/4,1/6,1/6)$ and m=5m=-5
  • $S^{2}(2,2,2,c)$, where $c=5,8,$ or $12$ and M=M(0;2,1/2,1/2,1/2,a/c)M=M(0;-2,1/2,1/2,1/2,a/c), $1<a<c/2$. In particular,
    • when $c=5$, $a=2$ and $m=3$
    • when $c=8$, $a=3$ and $m=3$
    • when $c=12$, $a=5$ and $m=5$
  • $S^{2}(2,2,3,4)$, where M=M(0;2,1/2,1/2,2/3,1/4)M=M(0;-2,1/2,1/2,2/3,1/4) and $m=5$
  • $S^{2}(2,2,4,6)$, where M=M(0;2,1/2,1/2,3/4,1/6)M=M(0;-2,1/2,1/2,3/4,1/6) and $m=7$
  • $S^{2}(2,2,2,2,c)$, where $c=4$ or $6$. In particular,
    • when $c=4$, M=M(0;2,1/2,1/2,1/2,1/2,1/4)M=M(0;-2,1/2,1/2,1/2,1/2,1/4) and m=3m=-3
    • when $c=6$, M=M(0;2,1/2,1/2,1/2,1/2,1/6)M=M(0;-2,1/2,1/2,1/2,1/2,1/6) and m=5m=-5
  • $S^{2}(2,2,2,2,2,2,2)$. See §5.3.
Lastly, we list the exceptions of Type (3). In each case, we list the only manifolds $M$, up to orientation, that admit a horizontal foliation with $e(\nu_{M})\neq 0$. One easily checks that in all cases $e(M)=0$:
  • $S^{2}(2,5,10)$, where M=M(0;1,1/2,1/5,3/10)M=M(0;-1,1/2,1/5,3/10) and M=M(0;1,1/2,2/5,1/10)M=M(0;-1,1/2,2/5,1/10)
  • $S^{2}(3,4,12)$, where M=M(0;1,1/3,1/4,5/12)M=M(0;-1,1/3,1/4,5/12)
  • $S^{2}(2,2,2,2,2,2)$, where M=M(0;3,1/2,1/2,1/2,1/2,1/2,1/2)M=M(0;-3,1/2,1/2,1/2,1/2,1/2,1/2)